OpenStudy (anonymous):

Evaluate for 5/6x for x = 6/7

OpenStudy (anonymous):

@narissa

OpenStudy (tkhunny):

Have you considered substituting the value given?

OpenStudy (anonymous):

what do you mean

OpenStudy (lochana):

\[ x = \frac{6}{7}\] \[ x\frac{5}{6} = \frac{6}{7}\frac{5}{6}\] \[ x\frac{5}{6} = \frac{5}{7}\]

OpenStudy (lochana):

is that what you are looking?

OpenStudy (anonymous):

yes thank you :p

OpenStudy (lochana):

bye